# Customized Interior-Point Methods Solver for Embedded Real-Time Convex Optimization

### v1

@@ III-E Solving the Linear System
[…]
$\tilde{K}=\left[\begin{array}[]{c|cc}Q&A^{T}&G^{T}\\ \hline\cr A&0&0\\ G&0&-W^{T}W\end{array}\right]+\left[\begin{array}[]{c|c}\delta_{s}I&0\\ \hline\cr 0&-\delta_{s}I\end{array}\right]$
@@
$P\tilde{K}P^{T}=LDL^{T}$
exists for any permutation $P$, and eliminates the need for data-dependent permutations, which are typically required in the $LDL^{T}$ factorization of general symmetric indefinite matrices [47]. As a result, the process of solving the linear system—which requires the most computational effort—can be formulated more efficiently.
@@
After obtaining the factors $L$ and $D$ through factorization, the linear system in (24) can be easily solved via backward substitution, diagonal scaling, and forward substitution [50, §3]. However, the resulting solution corresponds to the regularized system $\tilde{K}\Delta\xi_{i}=b_{i}$, rather than the original system $K\Delta\xi_{i}=b_{i}$. To recover the solution to the original system, we apply iterative refinement [51, §4.11] to progressively reduce the residual error $\|K\Delta\xi_{i}-b_{i}\|_{\infty}$ in $\Delta\xi_{i}$. Specifically, it iteratively solves the linear system
@@
$\Delta\xi_{i}\leftarrow\Delta\xi_{i}+\delta\xi_{i}$
until the residual error becomes sufficiently small. It is worth noting that no additional factorization is required to solve the linear system in (25), as it shares the same coefficient matrix $\tilde{K}$ with the system in (24). Therefore, only forward, backward substitution and diagonal scaling are needed during the iterative refinement process.
@@ III-F Initialization
$\begin{split}\underset{\hat{x},\hat{s}}{\text{minimize}}&\quad\frac{1}{2}\hat{x}^{T}Q\hat{x}+q^{T}\hat{x}+\frac{1}{2}\|\hat{s}\|_{2}^{2}\\ \text{subject to}&\quad G\hat{x}+\hat{s}=h,\quad A\hat{x}=b\end{split}$
Applying the optimality condition, the solution of the problem can be obtained by solving the following linear system:
$\begin{bmatrix}Q&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}\hat{x}\\ \hat{y}\\ \hat{z}\end{bmatrix}=\begin{bmatrix}-q\\ b\\ h\end{bmatrix}$
@@
$s^{(0)}=\begin{cases}-\hat{z}&\qquad(\alpha_{p}<0)\\ -\hat{z}+(1+\alpha_{p})e&\qquad(\text{otherwise})\end{cases}$
where $\alpha_{p}=\inf\{\alpha:-\hat{z}+\alpha e\in\mathcal{K}\}$ for $s^{(0)}$ to be located in the cone $\mathcal{K}$. Similarly for $z^{(0)}$, we set
$z^{(0)}=\begin{cases}\hat{z}&\qquad(\alpha_{d}<0)\\ \hat{z}+(1+\alpha_{d})e&\qquad(\text{otherwise})\end{cases}$
where $\alpha_{d}=\inf\{\alpha:\hat{z}+\alpha e\in\mathcal{K}\}$.
@@
which provides feasible $x$ with respect to the equality constraint $Ax=b$, and minimum norm of the residual for the inequality constraint. The solution of the problem (26) is obtained by solving the following linear system:
$\begin{bmatrix}0&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}\hat{x}\\ \hat{y}\\ -\hat{s}\end{bmatrix}=\begin{bmatrix}0\\ b\\ h\end{bmatrix}$
@@
$s^{(0)}=\begin{cases}\hat{s}&\qquad\alpha_{p}<0\\ \hat{s}+(1+\alpha_{p})e&\qquad\text{otherwise}\end{cases}$
where $\alpha_{p}=\inf\{\alpha:\hat{s}+\alpha e\in\mathcal{K}\}$. For initial dual variable $(y^{(0)},z^{(0)})$, we solve the constrained least-squares problem to minimize the norm of the inequality constraint in the dual problem.
@@
$\begin{bmatrix}0&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}x\\ \hat{y}\\ \hat{z}\end{bmatrix}=\begin{bmatrix}-q\\ 0\\ 0\end{bmatrix}$
@@
$z^{(0)}=\begin{cases}\hat{z}&\qquad\alpha_{d}<0\\ \hat{z}+(1+\alpha_{d})e&\qquad\text{otherwise}\end{cases}$
where $\alpha_{d}=\inf\{\alpha:\hat{z}+\alpha e\in\mathcal{K}\}$. We set $\kappa^{(0)}=1$, $\tau^{(0)}=1$ for both case of $Q=0$ and $Q\neq 0$.
@@ III-H Overall Algorithm
The overall algorithm of the proposed solver is summarized in Algorithm 1.
@@
0:  Tolerance $\epsilon_{\text{feas}}$, $\epsilon_{\text{gap}}$, $\epsilon_{\text{abs}}$ and $\epsilon_{\text{rel}}$ for termination criteria, and the problem data given as: $Q$, $q$, $A$, $b$, $G$, $h$ and $\mathcal{K}$.
@@
[…]
@@
[…]
@@
IV Solver Customization
This section presents the solver customization framework and the code generation tool developed in this study. A custom solver is tailored to handle a specific problem class that shares the same sparsity pattern in the data matrices $Q$, $A$, and $G$, as well as the same underlying structure of the cone $\mathcal{K}$. We refer to such a problem class as a problem family, and each individual problem within a problem family as a problem instance. The code generation tool allows users to easily obtain a custom solver for a desired problem family. The obtained custom solver is then integrated into the user application to solve problem instances belonging to the same problem family. The overall workflow for the customization and the code generation process is illustrated in Fig. 1.
@@ IV Solver Customization
IV-A Exploiting Sparsity Patterns of the KKT System
In this study, the sparsity pattern of the problem family is analyzed offline to perform AMD ordering and symbolic factorization which compute the sparsity pattern of the factor $L$ [50, §4]. Since these procedures depend solely on the sparsity pattern of $\tilde{K}$, rather than its numerical values, they can be computed once offline and applied to all problem instances within the same problem family. The resulting permutation and sparsity pattern of $L$ are cached in the custom solver code and utilized online during numerical factorization which computes the numerical values of the factors [50, §4]. This customization framework enables the custom solver to fully rely on static memory allocation of the problem data without the need for dynamic memory allocation. For numerical factorization, the algorithm proposed in [52] is implemented using nested for-loops, which enables code generation that is nearly independent of problem size [35, 36].
To enable efficient solver customization, we develop a code generation tool, implemented in MATLAB, that produces a custom solver code tailored to a given problem family. It analyzes the sparsity pattern of the problem and generates C code for a custom solver using a predefined code template. The generated code requires no external dependencies other than the standard library math.h for square root computations, and utilizes only the upper triangular part of the KKT matrix for its linear algebra operations, thereby reducing memory usage. The input interface of the code generation tool allows users to provide either the problem data for a specific problem instance or a symbolic description of the problem family using the MATLAB Symbolic Math Toolbox. Through the latter interface, the code generation tool produces not only the custom solver code but also the associated parsing information necessary to update the problem data for varying problem instances. (See bottom-left corner of Fig. 1 for an example of the parsing information provided by the code generation tool.) This feature particularly facilitates handling of problem data stored in compressed column storage (CCS) format, where direct access to individual elements is nontrivial. Furthermore, by providing symbolic expressions for the equations corresponding to each element of the problem data matrices, the tool significantly reduces the user’s workload when using the solver.
@@ V Benchmark Tests
In the benchmark tests, the performance of the solvers is evaluated using the performance profile [53], which provides a comprehensive comparison of solver performance across diverse problem collections. Let us consider the set of problem collections $\mathcal{P}$ and the set of solvers $\mathcal{S}$ for the current benchmark test. The relative performance ratio for solver $s\in\mathcal{S}$ and problem $p\in\mathcal{P}$ is defined as,
$r_{p,s}=\frac{t_{p,s}}{\displaystyle\min_{s\in\mathcal{S}}\left\{t_{p,s}\right\}}$
@@
$\rho^{r}_{s}(\tau)=\frac{1}{N_{\mathcal{P}}}\sum_{p\in\mathcal{P}}\mathcal{I}(r_{p,s};\tau)$
where $N_{\mathcal{P}}$ represents the number of problems in the problem set $\mathcal{P}$, and the indicator function $\mathcal{I}$ is defined as:
$\mathcal{I}(x;\tau)=\begin{cases}1&\text{if }x\leq\tau\\ 0&\text{otherwise}\end{cases}$
@@
$\rho_{s}^{a}(\tau)=\frac{1}{N_{\mathcal{P}}}\sum_{p\in\mathcal{P}}\mathcal{I}(t_{p,s};\tau)$
@@ V-B Portfolio Optimization
Portfolio optimization problem seeks to find an asset allocation that maximizes the risk-adjusted return [2, §4.7.6].
[…]
where $x\in\mathbb{R}^{n}$ represents the portfolio, $\mu\in\mathbb{R}^{n}$ is the vector of expected return, $\gamma>0$ is the risk-aversion parameter, and $\Sigma\in\mathbb{S}^{n}_{+}$ represents the risk model covariance matrix. This risk covariance model is typically expressed in factor form as:
$\Sigma=FF^{T}+D$
where $F\in\mathbb{R}^{n\times k}$ ($k<n$) is the factor-loading matrix, and $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix representing idiosyncratic risk. To simplify the problem, we introduce a new variable $y=F^{T}x$, which transforms the original formulation into:
$\begin{split}\underset{x,y}{\text{maximize}}&\quad\frac{1}{2}x^{T}Dx+\frac{1}{2}y^{T}y-\frac{1}{2\gamma}\mu^{T}x\\ \text{subject to}&\quad y=F^{T}x,\quad\mathbf{1}^{T}x=1\\ &\quad x\geq 0\end{split}$
[…]
Results—Figure 3 shows the benchmark results for portfolio optimization problems. We can observe that the solver developed in this study outperforms other solvers across all problem instances. These results demonstrate the effectiveness of the algorithmic formulation for handling quadratic cost functions, as well as the strength of IPMs in small- to medium-scale problems.
@@ V-C LASSO
[…]
$v_{i}\sim\begin{cases}0&\text{sampled with probability of 0.5}\\ \mathcal{N}(0,1/n)&\text{otherwise}\end{cases}$
Then, the vector $b$ can be constructed with noise $\epsilon_{i}\sim\mathcal{N}(0,1)$ as follows:
$b=Av+\epsilon$
Finally, the weighting parameter $\lambda$ is chosen as $(1/5)\|A^{T}b\|_{\infty}$. The total number of nonzero components in the problem data matrices ranges from 108 to 1,395,786.

### v2

@@ III-E Solving the Linear System
where i=1, 2i=1,\ 2. Since the KKT matrix $K$ is symmetric indefinite, we apply static regularization [55, 39] to render the KKT matrix quasi-definite, i.e., symmetric with the block located in the (1, 1) position being positive definite, the block in the (2, 2) position being negative definite, and the off-diagonal blocks being full-rank [55].
$\tilde{K}=\left[\begin{array}[]{c|cc}Q&A^{T}&G^{T}\\ \hline\cr A&0&0\\ G&0&-W^{T}W\end{array}\right]+\left[\begin{array}[]{c|c}\delta_{s}I&0\\ \hline\cr 0&-\delta_{s}I\end{array}\right]$ (45)
@@
$P\tilde{K}P^{T}=LDL^{T}$ (46)
exists for any permutation $P$, where $L$ denotes a lower triangular matrix and $D$ a diagonal matrix in the resulting factorization. This property eliminates the need for data-dependent permutations, which are typically required in the $LDL^{T}$ factorization of general symmetric indefinite matrices [11]. As a result, the process of solving the linear system—which requires the most computational effort—can be formulated more efficiently.
@@
After obtaining the factors $L$ and $D$ through factorization, the linear system in (44) can be easily solved via backward substitution, diagonal scaling, and forward substitution [16, §3]. However, the resulting solution corresponds to the regularized system $\tilde{K}\Delta\xi_{i}=b_{i}$, rather than the original system $K\Delta\xi_{i}=b_{i}$. To recover the solution to the original system, we apply iterative refinement [24, §4.11] to progressively reduce the residual error $\|K\Delta\xi_{i}-b_{i}\|_{\infty}$ in $\Delta\xi_{i}$. Specifically, it iteratively solves the linear system
@@
$\Delta\xi_{i}\leftarrow\Delta\xi_{i}+\delta\xi_{i}$ (48)
until $\|K\Delta\xi_{i}-b_{i}\|_{\infty}<\epsilon_{ir}$, where the default tolerance is set to $\epsilon_{ir}=10^{-13}$. It is worth noting that no additional factorization is required to solve the linear system in (47), as it shares the same coefficient matrix $\tilde{K}$ with the system in (44). Therefore, only forward, backward substitution and diagonal scaling are needed during the iterative refinement process.
@@ III-F Initialization
$\begin{split}\underset{\hat{x},\hat{s}}{\text{minimize}}&\quad\frac{1}{2}\hat{x}^{T}Q\hat{x}+q^{T}\hat{x}+\frac{1}{2}\|\hat{s}\|_{2}^{2}\\ \text{subject to}&\quad G\hat{x}+\hat{s}=h,\quad A\hat{x}=b\end{split}$ (49)
where $\hat{(\cdot)}$ denotes the auxiliary variables associated with the initialization problem. Applying the optimality condition, the solution of the problem can be obtained by solving the following linear system:
$\begin{bmatrix}Q&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}\hat{x}\\ \hat{y}\\ \hat{z}\end{bmatrix}=\begin{bmatrix}-q\\ b\\ h\end{bmatrix}$ (50)
@@
$s^{(0)}=\begin{cases}-\hat{z}&\qquad\alpha_{p}<0\\ -\hat{z}+(1+\alpha_{p})e&\qquad\text{otherwise}\end{cases}$ (51)
where $\alpha_{p}=\inf\{\alpha:-\hat{z}+\alpha e\in\text{int}\mathcal{K}\}$ for $s^{(0)}$ to be located in the interior of the cone $\mathcal{K}$. Similarly for $z^{(0)}$, we set
$z^{(0)}=\begin{cases}\hat{z}&\qquad\alpha_{d}<0\\ \hat{z}+(1+\alpha_{d})e&\qquad\text{otherwise}\end{cases}$ (52)
where $\alpha_{d}=\inf\{\alpha:\hat{z}+\alpha e\in\text{int}\mathcal{K}\}$.
@@
which provides feasible $x^{(0)}$ with respect to the equality constraints, and minimum norm of the residual for the inequality constraints. The solution of the problem (53) is obtained by solving the following linear system:
$\begin{bmatrix}0&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}\hat{x}\\ \hat{y}\\ -\hat{s}\end{bmatrix}=\begin{bmatrix}0\\ b\\ h\end{bmatrix}$ (54)
@@
$s^{(0)}=\begin{cases}\hat{s}&\qquad\alpha_{p}<0\\ \hat{s}+(1+\alpha_{p})e&\qquad\text{otherwise}\end{cases}$ (55)
where $\alpha_{p}=\inf\{\alpha:\hat{s}+\alpha e\in\text{int}\mathcal{K}\}$. For initial dual variable $(y^{(0)},z^{(0)})$, we solve the constrained least-squares problem to minimize the norm of the residual for the inequality constraints in the dual problem.
@@
$\begin{bmatrix}0&A^{T}&G^{T}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}\hat{x}\\ \hat{y}\\ \hat{z}\end{bmatrix}=\begin{bmatrix}-q\\ 0\\ 0\end{bmatrix}$ (57)
@@
$z^{(0)}=\begin{cases}\hat{z}&\qquad\alpha_{d}<0\\ \hat{z}+(1+\alpha_{d})e&\qquad\text{otherwise}\end{cases}$ (58)
where $\alpha_{d}=\inf\{\alpha:\hat{z}+\alpha e\in\text{int}\mathcal{K}\}$. We set $\kappa^{(0)}=1$, $\tau^{(0)}=1$ for both case of $Q=0$ and $Q\neq 0$.
@@ III-H Overall Algorithm
The overall algorithm of the proposed solver is summarized in Algorithm 1. Here, the variable FLAG denotes the solution status indicator, where the enum-type status values OPT, PINF, and DINF represent successful convergence to an optimal solution, primal infeasibility, and dual infeasibility, respectively.
IV Solver Customization
This section presents the solver customization framework and the code generation tool developed in this study. A custom solver is tailored to handle a specific problem class that shares the same sparsity pattern in the data matrices $Q$, AA, and $G$, as well as the same underlying structure of the cone $\mathcal{K}$. We refer to such a problem class as a problem family, and each individual problem within a problem family as a problem instance. The code generation tool allows users to easily obtain a custom solver for a desired problem family. The obtained custom solver is then integrated into the user application to solve problem instances belonging to the same problem family. The overall workflow for the customization and the code generation process is illustrated in Fig. 1.
@@ IV Solver Customization
0: Tolerance $\epsilon_{\text{feas}}$, $\epsilon_{\text{gap}}$, $\epsilon_{\text{abs}}$ and $\epsilon_{\text{rel}}$ for termination criteria, and the problem data given as: $Q$, $q$, AA, $b$, $G$, $h$ and $\mathcal{K}$.
@@
8:    $\texttt{FLAG}\leftarrow\texttt{PINF}$ and break.
@@
10:    $\texttt{FLAG}\leftarrow\texttt{DINF}$ and break.
@@
IV-A Exploiting Sparsity Patterns of the KKT System
In this study, the sparsity pattern of the problem family is analyzed offline to perform AMD ordering and symbolic factorization which compute the sparsity pattern of the factor $L$ [16, §4]. Since these procedures depend solely on the sparsity pattern of $\tilde{K}$, rather than its numerical values, they can be computed once offline and applied to all problem instances within the same problem family. The resulting permutation and sparsity pattern of $L$ are cached in the custom solver code and utilized online during numerical factorization which computes the numerical values of the factors [16, §4]. This customization framework enables the custom solver to fully rely on static memory allocation of the problem data without the need for dynamic memory allocation (i.e., no use of memory allocation functions such as malloc during implementation). For numerical factorization, the algorithm proposed in [15] is implemented using nested for-loops, which enables code generation that is nearly independent of problem size [21, 5].
To enable efficient solver customization, we develop a code generation tool, implemented in MATLAB, that produces a custom solver code tailored to a given problem family. It analyzes the sparsity pattern of the problem and generates C code for a custom solver using a predefined code template. The generated code requires no external dependencies other than the standard library math.h for square root computations, and utilizes only the upper triangular part of the KKT matrix for its linear algebra operations, thereby reducing memory usage compared to using the full matrix. The input interface of the code generation tool allows users to provide either the problem data for a specific problem instance or a symbolic description of the problem family using the MATLAB Symbolic Math Toolbox. Through the latter interface, the code generation tool produces not only the custom solver code but also the associated parsing information necessary to update the problem data for varying problem instances. (See bottom-left corner of Fig. 1 for an example of the parsing information provided by the code generation tool.) This feature particularly facilitates handling of problem data stored in compressed column storage (CCS) format, where direct access to individual elements is nontrivial. Furthermore, by providing symbolic expressions for the equations corresponding to each element of the problem data matrices, the tool significantly reduces the user’s workload when using the solver.
@@ V Benchmark Tests
In the benchmark tests, the performance of the solvers is evaluated using the performance profile [19], which provides a comprehensive comparison of solver performance across diverse problem collections. Let us consider the set of problem collections $\mathbf{P}$ and the set of solvers $\mathbf{S}$ for the current benchmark test. The relative performance ratio for solver $s\in\mathbf{S}$ and problem $p\in\mathbf{P}$ is defined as,
$r_{p,s}=\frac{t_{p,s}}{\displaystyle\min_{s\in\mathbf{S}}\left\{t_{p,s}\right\}}$ (62)
@@
$\rho^{r}_{s}(\tau)=\frac{1}{N_{\mathbf{P}}}\sum_{p\in\mathbf{P}}\mathcal{I}(r_{p,s};\tau)$ (63)
where $N_{\mathbf{P}}$ represents the number of problems in the problem set $\mathbf{P}$, and the indicator function $\mathcal{I}$ is defined as:
$\mathcal{I}(x;\tau)=\begin{cases}1&\text{if }x\leq\tau\\ 0&\text{otherwise}\end{cases}$ (64)
@@
$\rho_{s}^{a}(\tau)=\frac{1}{N_{\mathbf{P}}}\sum_{p\in\mathbf{P}}\mathcal{I}(t_{p,s};\tau)$ (65)
@@ V-B Portfolio Optimization
Portfolio optimization problem seeks to find an asset allocation that maximizes the risk-adjusted return [9, §4.7.6].
$\begin{split}\underset{\omega}{\text{maximize}}&\quad\bar{\mu}^{T}\omega-\rho\left(\omega^{T}\Sigma\omega\right)\\ \text{subject to}&\quad\mathbf{1}^{T}\omega=1\\ &\quad\omega\geq 0\end{split}$ (66)
where $\omega\in\mathbb{R}^{n}$ represents the portfolio, $\bar{\mu}\in\mathbb{R}^{n}$ is the vector of expected return, $\rho>0$ is the risk-aversion parameter, and $\Sigma\in\mathbb{S}^{n}_{+}$ represents the risk model covariance matrix. This risk covariance model is typically expressed in factor form as:
$\Sigma=FF^{T}+\Pi$ (67)
where $F\in\mathbb{R}^{n\times k}$ ($k<n$) is the factor-loading matrix, and $\Pi\in\mathbb{R}^{n\times n}$ is a diagonal matrix representing idiosyncratic risk. To simplify the problem, we introduce a new variable $\zeta=F^{T}\omega$, which transforms the original formulation into:
$\begin{split}\underset{\omega,\zeta}{\text{maximize}}&\quad\frac{1}{2}\omega^{T}\Pi\omega+\frac{1}{2}\zeta^{T}\zeta-\frac{1}{2\rho}\bar{\mu}^{T}\omega\\ \text{subject to}&\quad\zeta=F^{T}\omega,\quad\mathbf{1}^{T}\omega=1\\ &\quad\omega\geq 0\end{split}$ (68)
making the Hessian of the quadratic cost function diagonal. The problem instances for benchmark test are generated following the similar approach used in [51]. The instances are generated by increasing number of factors k=[5, 10, 20, 50, 100]k=[5,\ 10,\ 20,\ 50,\ 100] and ratios of n/k=[10, 20, 50, 100]n/k=[10,\ 20,\ 50,\ 100]. For each problem size, five different problems are generated, thus total 100 instances are used for the benchmark test. The elements of the $F$ matrix are sampled as $F_{i,j}\sim\mathcal{N}(0,1)$ with 70% nonzero elements, the elements of $\Pi$ matrix are sampled as $\Pi_{i,i}\sim\mathcal{U}[0,\sqrt{k}]$, and mean return $\bar{\mu}$ is sampled as $\bar{\mu}_{i}\sim\mathcal{N}(0,1)$. The total number of nonzero components in the problem data matrices ranges from 338 to 54,355.
Results—Figure 3 shows the benchmark results for portfolio optimization problems. We can observe that the solver developed in this study outperforms the other solvers across all problem instances. The problem instances in this benchmark set correspond to small- to medium-scale problems, where IPM-based solvers generally exhibit better performance than the ADMM-based solver. Moreover, among the IPM-based solvers, the developed solver achieves higher computational efficiency owing to its ability to directly handle quadratic objective functions. This capability allows matrix factorization to be performed on a smaller and sparser KKT matrix than those of the other IPM-based solvers, ECOS and MOSEK, which are limited to linear objective functions. Therefore, these results demonstrate the effectiveness of the algorithmic formulation for handling quadratic cost functions, as well as the strength of IPMs in small- to medium-scale problems.
@@ V-C LASSO
$\begin{split}\underset{\theta}{\text{minimize}}&\qquad\frac{1}{2}\|X\theta-\bar{y}\|_{2}^{2}+\bar{\lambda}\|\theta\|_{1}\end{split}$ (69)
where $\theta\in\mathbb{R}^{n}$ is the vector of parameters, $X\in\mathbb{R}^{m\times n}$ is the data matrix, and $\bar{\lambda}>0$ is the weighting parameter. For LASSO, the problem instances for benchmark test are also generated following the method used in [51]. We increase the number of regression parameter n=[10, 50, 100, 500, 700]n=[10,\ 50,\ 100,\ 500,\ 700] and the ratio m/n=[2, 5, 10,20]m/n=[2,\ 5,\ 10,20]. For each problem size, five different problems are generated, thus total 100 instances are used for the benchmark test. The elements of matrix $X$ are sampled as $X_{i,j}\sim\mathcal{N}(0,1)$ with 15% nonzero elements. To construct the vector $\bar{y}$, we construct the true sparse vector $v\in\mathbb{R}^{n}$ to be learned as:
$v_{i}\sim\begin{cases}0&\text{sampled with probability of 0.5}\\ \mathcal{N}(0,1/n)&\text{otherwise}\end{cases}$ (70)
Then, the vector $\bar{y}$ can be constructed with noise $\epsilon_{i}\sim\mathcal{N}(0,1)$ as follows:
$\bar{y}_{i}=\sum_{j=1}^{n}X_{i,j}v_{j}+\epsilon_{i}\quad(i=1,2,\cdots,m)$ (71)
Finally, the weighting parameter $\bar{\lambda}$ is chosen as $(1/5)\|X^{T}\bar{y}\|_{\infty}$. The total number of nonzero components in the problem data matrices ranges from 108 to 1,395,786.